A point source of light is 6 ft. from the center of a sphere with a radius of 3 ft. Find surface area that is brightened.


1
Expert's answer
2021-09-27T01:44:08-0400

"S.A=4πr^2\\
3^2=9\\
=4\cdot \:9\pi \\
=36\pi"